What is the t value with a 90% confidence interval for the true population mean if the sample size n = 27?

(Please keep three decimal places)

t value= ____________


1
Expert's answer
2022-05-10T23:03:46-0400

The critical value for "\alpha = 0.1" and "df = n-1 = 26" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 1.706."